OpenStudy (ujjwal):

If the two normals at \(t_1\) and \(t_2\) on parabola \(y^2=4ax\) meet the parabola again at \(t_3\) then \(t_1t_2=\): a)1 b)-1 c)2 d)-2

OpenStudy (experimentx):

t1 and t2 are values of x or y?

OpenStudy (ujjwal):

I guess something parametric.. Not very sure.. It has something to do with parametric equations!

OpenStudy (turingtest):

usually you are given formulas for x and y in terms of t in these situations. are you sure you haven't omitted any info from the question?

OpenStudy (ujjwal):

I am 100% sure i haven't omitted anything..

OpenStudy (anonymous):

parametric eqs for y^2 = 4ax will be x = 2at and y = at^2.....this is the standard form.

OpenStudy (ujjwal):

exactly! but i don't know how to proceed.. :(

OpenStudy (anonymous):

oh I'm sorry....I gave the parametric eqs wrong...it should be x = at^2 and y = 2at.

OpenStudy (ujjwal):

Did you?? LOL.. I didn't even read them.. I am reading your answer right now.

OpenStudy (anonymous):

yea...my bad! just realized i wrote them the other way round!

OpenStudy (experimentx):

let y=t, x=t^2/4a, ...  perhaps